Jezreel borrowed PHP250,000.00 from a bank at 6% discount rate. How much would be receive from the bank if he agreed to pay the amount after 5 years.
Answer by Theo(13342)

he borrowed 250,000 from the bank at 6% discount rate.

he agreed to pay them back that amount in 5 years.

what i understand from this is that the bank expects him to pay them 250,000 in 5 years.

since the bank needs to make 6% a year on their investment, you are looking for the present value of a future value of 250,000.

the formula to use is f = p * (1 + r) ^ n

f is the future value
p is the present value
r is the interest rate per time period.
n is the number of time periods.

in this problem, the formula becomes:

250,000 = p * 1.06 ^ 5.

divide both sides of this equation by 1.06 ^ 5 to get:

250,000 / 1.06 ^ 5 = p

solve for p to get:

p = 186,814.5432.

that's how much he would get now if he promises to pay the bank 250,000 five years from now.

if he needed 250,000 now, then the formula becomes:

f = 250,000 * 1.06 ^ 5 = 334,556.3944.

that's how much he would have to pay the bank back in 5 years if he needed to have 250,000 now.